# DocNet: Semantic Structure in Inductive Bias Detection Models

### v1

@@ 4 Results and Discussion
We ran 1284 experiments to train multiple embeddings and supervised learning models over the various configurations and four datasets, to include a combined topic agnostic dataset that merges the AFG, OATH, and VAX datasets. The experiments are trained on 1 node of an NVIDIA DGX A100 system. Each graph autoencoder embedding model takes at most 2 GPU hours to train one model, while each Graph2Vec embedding model takes at most 3.5 CPU minutes. In order to control for different configurations, we formally evaluate the difference between configurations and embeddings using a one-sided Wilcoxan signed rank test. This allows us to determine the paired difference between two methods when all other configurations are held constant. We are able to confirm that the embeddings are valuable in comparison to the naive baseline for all metrics, with p-values close to 0. Our experiments have an average of .21 increase in macro F-1 scores over the baseline. This is on average a 120% improvement over the baseline.
Table 3: Average (std) of results across all embedding types
@@
graph2vec_no_features 0.498528 (.169) 0.341580 (.137)
GPT 3.5 Turbo 0.416704 (.164) 0.349551 (.177)
VGAEautoencode_data_sent 0.474973 (.143) 0.357707 (.128)
VGAEautoencode_data 0.470229 (.155) 0.362188 (.135)
Llama 3-8B 0.453731 (.225) 0.378090 (.204)
GAEautoencode_data_sent 0.524193 (.146) 0.391000 (.135)
GAEautoencode_data 0.537793 (.146) 0.425286 (.138)
doc2vec_dbow_graph2vec_no_features 0.559085 (.169) 0.456914 (.170)
doc2vec_dm_graph2vec_no_features 0.543820 (.178) 0.466900 (.184)
doc2vec_dm 0.538900 (.170) 0.472554 (.170)
doc2vec_dbow 0.573575 (.170) 0.485918 (.170)
doc2vec_dm_graph2vec_nodewords 0.583128 (.164) 0.500872 (.167)
graph2vec_spacy 0.588646 (.154) 0.501290 (.161)
doc2vec_dbow_graph2vec_spacy 0.584941 (.163) 0.501290 (.165)
graph2vec_nodewords 0.591630 (.151) 0.506795 (.157)
doc2vec_dm_graph2vec_spacy 0.572459 (.157) 0.508879 (.160)
doc2vec_dm_graph2vec_nodewords 0.570328 (.164) 0.510572 (.167)
SBERT 0.606528 (.154) 0.541717 (.162)
@@
The results of models grouped on different label bins are in Table 4. The large variation in scores is primarily a result of the increased difficulty with the greater number of classes. We find that after regrouping the “Full” label models’ predictions, there is not a statistically significant difference between models trained on the “Full” label set in comparison to “LCR” or “Binary” labels, p-value of .91 and .36, respectively. This demonstrates that while the “Full” multiclass models have inherently lower scores, they are able to extract divisions in the news’ embedding space just as successfully as the models trained on fewer categories.
Table 4: Average (std) of results based on different label aggregations
@@
Full 0.403810 (.102) 0.293628 (.111) .379535 (.075) .105336 (.038)
LCR 0.517936 (.102) 0.465078 (.102) .470023 (.067) .216987 (.036)
Binary 0.722017 (.099) 0.608042 (.118) .7166878 (.069) .416615 (.024)
@@
The results for the different domain aggregation comparisons are listed in Table 5. We find that with the BASIL dataset, the embeddings separated more easily into domain level labels (although the article models were still significantly better than the naive baseline across all metrics with a p-value near 0). We also find that the method of aggregating article embeddings makes a significant difference on performance, such that domain models using the Frobenius norm of the difference from its topic perform worse than news level predictions. The mean of article embeddings across all samples of a domain has the best performance in domain level bias detection.
Table 5: Average (std) of results across additional configurations
@@
froNormTopicDiff 0.500561 (.159) 0.397489 (.153)
no_aggregation 0.564704 (.165) 0.453129 (.167)
avgTopicDiff 0.555718 (.163) 0.483041 (.167)
arithmeticMean 0.563010 (.168) 0.489796 (.175)
BASIL_articles 0.372284 (.111) 0.332773 (.112)
BASIL_domains 0.587901 (.190) 0.543800 (.196)
@@
no_filter 0.520002 (.173) 0.445451 (.174))
filterAfg 0.549137 (.157) 0.447124 (.143)
filterVaccine 0.510295 (.148) 0.450861 (.166)
filterMultArticles 0.535061 (.168) 0.452376 (.168)
filterMultTopic 0.578134 (.171) 0.463696 (.195)
4.1 Embedding Model
@@
The average mean and standard deviation of scores given different embeddings and models are in Table 3. Through the paired rank test on the full set of experiments, we find that SBERT outperforms all other embeddings using a p-value significance threshold of .05. The GPT, Llama, graph autoencoders, and Graph2Vec with no features (i.e. node attributes as degrees) perform the worst, with no statistically significant performance difference between them. The unsupervised Graph2Vec based embeddings are able to capture characteristics of biased text with performance often times similar to, and in the case of “graph2vec_nodewords”, 30% of the time, better than SBERT. Graph2Vec with node attributes (both with just node words and word vectors) outperforms both doc2vec models with p-values all less than .01. We also observe that when Graph2Vec embeddings with node features are combined with doc2vec, a statistically significant improvement (p-values are all less than .01) from the isolated doc2vec model occurs. For instance, Graph2Vec with node words concatenated with doc2vec with distributed memory (“doc2vec_dm_graph2vec_nodewords”) consistently outperforms embeddings using just “doc2vec_dm”, as observed in the their average scores in Table 3. There is no statistical difference between the type of node attribute used in Graph2Vec (node words or Spacy vectors).
@@
These results demonstrate the value of including semantic structure in bias classification so long as there is a unique node identifier that can be used to indirectly relate separate graphs. Our Graph2Vec models do not factor in edge attributes, and its only information on word frequency is the lack of certain edges due to non positive PMI. Unlike with SBERT, language models and word embeddings also are not necessary in this characterization. An indirect tie between document networks is necessary and sufficient to have performance that often times is comparable to transformer based document embeddings. Therefore, characterizing text using individual co-occurrence graphs results in improved bias detection.
@@
While our computationally expensive graph autoencoder models pass both edge and node vector attributes in generating document embeddings, they still under perform in comparison to the simple Graph2Vec unsupervised embeddings. Further experimentation and tuning should be conducted to validate if alternative deep learning graph models, like graph attention or diffusion based approaches Nikolentzos et al. (2020); Liu et al. (2021), could be repurposed for inductive embedding generation, as graph neural network based approaches are often prone to oversmoothing. Regardless, the fact that a simple approximate graph isomorphism comparison through WL-hashing with minimal additional attributes can generate good bias predictions, validates the value of characterizing semantic structure using document networks.
4.2 DocNet Characteristic Case Study
We finally explore if the models that use Graph2Vec embeddings are aligning along trivial graph features. In Figure 1 we plot graphs from the BASIL dataset using a spring force-directed algorithm [^3] https://networkx.org/documentation/stable/reference/generated/networkx.drawing.layout.spring_layout.html. We select articles that have the highest confidence (i.e. probability) predictions for the respective incorrect or correct binary bias labels from a Graph2Vec with node words embedding. We observe that in Figure 1, the graphs for the articles that were predicted to be unbiased (a and d) both have visually sparse layouts. Meanwhile, the biased articles (b and c) appear to be denser. Upon returning to the original data, we also find that the news article from Figure 1d is actually labeled at the article level to be unbiased. However, the original article label for Figure 1c is unbiased. This article  [^4] https://archive.nytimes.com/www.nytimes.com/politics/first-draft/2016/02/25/2-ex-presidents-of-mexico-say-no-way-theyre-.paying-for-donald-trumps-wall contains phrases like “Donald J. Trump leads a familiar refrain” and “Mr. Trump, who has been criticized for lacing his campaign speeches with profanity”. This causes us to question the validity of this article’s human annotated bias label. Furthermore, these examples demonstrate the continued difficulty in developing good labels for an inherently subjective application. Yet these graph plots do imply that there is a pattern in the graph structure that captures bias.
@@ 4.2 DocNet Characteristic Case Study
(c) Incorrect-Pred Biased
(d) Incorrect-Pred Unbiased
Figure 1: Graphs of articles with high probability predictions from BASIL. Predictions from graph2vec_nodewords at the domain level and binary labels (Macro F-1 = .82). Source domains: a) NYTimes, b) Fox, c) NYTimes d) Huffington Post
To further explore the characteristics of the semantic structure, we obtain various descriptive graph metrics for each article and calculate their correlation with bias labels as shown in Figures 2 and 3. In both correlation plots, there are no strong correlations between any of the metrics and the corresponding bias labels. We then train a simple logistic regression model on these statistics for each dataset to predict binary bias labels. These models have a .21 lower Macro F-1 on the partisan dataset and a .12 lower Macro F-1 on the domain level Basil predictions compared to the corresponding Graph2Vec with node words model. This analysis further demonstrates that while the semantic structure holds bias information, there is additional information to be gained by using our proposed graph embeddings rather than curated descriptive graph metrics.
@@ 5 Conclusion and Future Work
Through our graph embeddings we demonstrate that encoding semantic structure in news articles as co-occurrence networks can be used for both domain and article level bias detection. We couple baseline embeddings with a logistic regression to learn a separation of the embedding space and learn that bias texts have similar word context agnostic semantic structures that can be captured by document networks (i.e. DocNet). This is demonstrated by the success of the Graph2Vec embeddings in the supervised models, especially as it often has comparable performance to the SBERT embeddings, even without the use of word embeddings. DocNet does not require pre-trained language models to generate word embeddings, nor does it require computationally expensive deep learning techniques. The most computationally time consuming portion of DocNet is the initial graph construction of an article. However, this can be easily parallelized across a corpus of documents. It is completely inductive, allowing for inexpensive prediction on new datasets.
Through the various groupings of bias labels, we also identify that there is often insufficient information in the semantic structure of an article for accurate fine-grained bias classification, although future research using one-vs-rest classifiers rather than our generalized multinomial regression may be able to improve upon our results. This is an observation observed by other recent work Carragher et al. (2024). On the other hand, fuzzier binary classifications can still be sufficiently characterized by semantic structures. This means that right-leaning news articles and left-leaning news articles have significantly more in common than either side of the political spectrum may believe. Although we are not able to pinpoint exactly what this writing style bias is due to the lack of interpretability in our graph embeddings, we can definitively conclude that the characteristics of partisan rhetoric are more than context-dependent.
In future work, it would be valuable to research embeddings that can fuse word embeddings with semantic structure. While our graph autoencoders performed better than a naive baseline, they were significantly less useful than our Graph2Vec embeddings in bias detection. Context represented in pre-trained language models is valuable to bias detection, as demonstrated in our SBERT model. Future developments in inductive graph embedding approaches that can efficiently and effectively represent dense networks and attributes would improve bias detection efforts. In the interim, our DocNet methodology provides a pathway for low-resourced bias detection, where computational resources, rigorously labeled datasets, or language models are not available.

### v2

@@ 4 Results and Discussion
We ran 1284 experiments to train multiple embeddings and supervised learning models over the various configurations and four datasets, to include a topic agnostic dataset that merges the AFG, OATH, and VAX datasets, separate from BASIL. The experiments are trained on 1 node of an NVIDIA DGX A100 system. Each graph autoencoder embedding model takes at most 2 GPU hours to train one model, while each Graph2Vec embedding model takes at most 3.5 CPU minutes. In order to control for different configurations, we formally evaluate the difference between configurations and embeddings using a one-sided Wilcoxan signed rank test on the accuracy and macro F-1 scores. As each pairwise configuration uses the same test sets but are trained independently, the necessary assumptions for a valid hypothesis test are satisfied. This test allows us to determine if there is a paired difference between two methods when all other configurations are held constant.
Through this hypothesis test, We are able to confirm that our graph based embeddings are valuable in comparison to the naive baseline for both metrics, with p-values close to 0. Our experiments have an average .21 increase in macro F-1 scores over the baseline. This is on average a 120% improvement. For the sake of brevity, we display the average and standard deviation of scores for specific configurations in the following tables.
@@
naive baseline 0.522 (.159) 0.246 (.133)
graph2vec_no_features 0.498(.169) 0.341 (.137)
GPT 3.5 Turbo 0.417 (.164) 0.349 (.177)
VGAEautoencode_sent 0.475 (.143) 0.358 (.128)
VGAEautoencode 0.470 (.155) 0.362 (.135)
Llama 3-8B 0.454 (.225) 0.378 (.204)
GAEautoencode_sent 0.524 (.146) 0.391 (.135)
GAEautoencode 0.538 (.146) 0.425 (.138)
doc2vec_dbow_graph2vec_no_features 0.559 (.169) 0.457 (.170)
doc2vec_dm_graph2vec_no_features 0.544 (.178) 0.467 (.184)
doc2vec_dm 0.539 (.170) 0.472 (.170)
doc2vec_dbow 0.574 (.170) 0.486 (.170)
doc2vec_dm_graph2vec_nodewords 0.583 (.164) 0.501 (.167)
graph2vec_spacy 0.589 (.154) 0.501 (.161)
doc2vec_dbow_graph2vec_spacy 0.585 (.163) 0.501 (.165)
graph2vec_nodewords 0.592 (.151) 0.507 (.157)
doc2vec_dm_graph2vec_spacy 0.572 (.157) 0.509 (.160)
doc2vec_dm_graph2vec_nodewords 0.570 (.164) 0.511 (.167)
SBERT 0.607 (.154) 0.542 (.162)
Table 3: Accuracy and Macro F-1 given various embedding models.
Avg (std) of all data configurations
@@
We identify that, as expected, there is a large drop in performance as the number of classes increase, with the models trained for fine grained bias classification contributing to the lower overall average scores. The results from models grouped by various label bins are in Table 4. Of note, we find that after regrouping the predictions from models trained using “Full” labels, there is not a statistically significant difference between models trained on the “Full” label set in comparison to “LCR” or “Binary” labels, with p-values of .91 and .36, respectively. This demonstrates that while the “Full” multiclass models have low scores, they are able to extract characteristics of bias in the news’ embedding space just as successfully as the models trained on fewer categories. A one-vs rest model training approach may be used to validate if more granular bias detection is possible with these embeddings.
@@
Full 0.404 (.102) 0.294 (.111) .380 (.075) .105 (.038)
LCR 0.518 (.102) 0.465 (.102) .470 (.067) .216987 (.036)
Binary 0.722 (.099) 0.608 (.118) .717 (.069) .417 (.024)
Table 4: Accuracy and Macro F-1 given various label aggregation methods.
Avg (std) of all embedding models. The baseline scores are from the naive zero rule baseline
@@
The results for the different domain aggregation comparisons are listed in Table 5. We find that with the BASIL dataset, the embeddings separated more easily into domain level labels (although the article models were still significantly better than the naive baseline across all metrics with a p-value near 0). We also find that the method of aggregating article embeddings makes a significant difference on performance, such that domain models using the Frobenius norm of the difference from its topic perform the worst. The mean of article embeddings across all samples of a domain has the best performance in domain level bias detection. This could both imply that topic information is not significant in domain level bias detection or that a better way of extracting topic level information is warranted.
@@
froNormTopicDiff 0.500 (.159) 0.397 (.153)
no_aggregation 0.565 (.165) 0.453 (.167)
avgTopicDiff 0.556 (.163) 0.483 (.167)
arithmeticMean 0.563 (.168) 0.490 (.175)
BASIL_articles 0.372 (.111) 0.332 (.112)
BASIL_domains 0.588 (.190) 0.544 (.196)
@@
BASIL 0.480 (.189) 0.438 (.191)
no filter 0.520 (.173) 0.445 (.174)
AFG 0.549 (.157) 0.447 (.143)
VAX 0.510 (.148) 0.451 (.166)
OATH 0.560 (.161) 0.461 (.160)
Table 5: Accuracy and Macro F-1 given additional configurations of domain aggregation and article subsets. Average (std)
Embedding Model
@@
The average mean and standard deviation of scores given different embedding models are in Table 3. Through the paired rank test on the full set of experiments, the GPT, Llama, graph autoencoders, and Graph2Vec with no features (i.e. using the base graph) perform the worst, with no statistically significant performance difference between them. The unsupervised Graph2Vec based embeddings with node attributes are able to capture characteristics of biased text with performance often times similar to, and in the case of “graph2vec_nodewords”, 30% of the time, better than SBERT. However, we find that on average SBERT still outperforms all other embeddings using a p-value significance threshold of .05. Graph2Vec with node attributes (both with node words and word vectors) outperforms all Doc2Vec embedding models with p-values all less than .01. We also observe that when Graph2Vec embeddings with node features are combined with Doc2Vec, a statistically significant improvement (p-values are all less than .01) from the isolated Doc2Vec model occurs. For instance, Graph2Vec with node words concatenated with Doc2Vec with distributed memory (“doc2vec_dm_graph2vec_nodewords”) consistently outperforms embeddings using just Doc2Vec, as observed in their average scores in Table 3. We are not able to identify a statistical difference between embedding performance based on the type of node attribute used in Graph2Vec (node words or Spacy vectors).
@@
Although our graph based embeddings do not consistently outperform SBERT, these results still demonstrate the value of including semantic structure in bias classification. Unlike with SBERT, language models and word embeddings also are unnecessary in this characterization. As demonstrated by the performance of “graph2vec_nodewords”, so long as there is a unique node identifier that can be used to indirectly relate separate graphs, the article graph embeddings are comparably valuable to SBERT. This indirect connection between document networks is necessary and sufficient to have performance that comparable to language model based document embeddings. Without factoring in edge attributes, pre-trained word embeddings, or word frequency outside of missing edges due to non positive PMI, our document networks are able to successfully embed information for bias detection.
@@
On the other hand, even as our computationally expensive graph autoencoder models pass both edge and node attributes in generating document embeddings, they still under perform in comparison to the Graph2Vec unsupervised embeddings. Further experimentation and tuning could be conducted to validate if alternative deep learning graph models, like graph attention or diffusion based approaches (Nikolentzos, Tixier, and Vazirgiannis 2020; Liu et al. 2021), could be repurposed for inductive embedding generation, since graph neural network based approaches are often prone to oversmoothing, or if new graph hashing algorithms that efficiently encode edge attributes could be used in Graph2Vec. Regardless, the fact that a simple approximate graph isomorphism comparison through WL-hashing using minimal additional attributes can generate good bias predictions, validates the value of characterizing semantic structure using document networks.
DocNet Characteristics Case Study
In the following section, we investigate if the models that use Graph2Vec embeddings are extracting distinct graph features that align with bias labels.
To further explore the characteristics of the semantic structure, we obtain descriptive graph metrics for every article and calculate their Pearson correlation with an ordinal encoding of the bias labels as shown in Figures 1 and 2. In both correlation plots, there is not a strong correlation between any of the metrics and the corresponding bias labels. We then train a logistic regression model for each dataset to predict binary bias labels using these metrics. These models have an average .21 lower Macro F-1 on the topic aligned datasets and a .12 lower Macro F-1 on the domain level Basil predictions compared to the corresponding Graph2Vec with node words models. This analysis demonstrates that while the semantic structure of articles holds bias information, there is additional information to be gained by using our proposed graph embeddings rather than curated descriptive graph metrics.
We also plot graphs from the top performing binary graph2vec_node_words embedding models to manually inspect for any consistent characteristics or discrepancies in the data labels. We highlight graphs from the domain label BASIL model, and no domain aggregation VAX, and OATH datasets in Figures 3, 4 and 5, respectively. These graphs are the articles with the highest probability incorrect and correct predictions for both bias labels. The document networks are graphed using a spring force-directed algorithm (Hagberg, Swart, and S Chult 2008).
@@
(c) Incorrect-True Biased
(d) Incorrect-True Unbiased
Figure 3: Graphs of articles with high probability predictions from BASIL. Predictions from graph2vec_nodewords at the domain level and binary labels (Macro F-1 = .82). Source domains: a) NYTimes, b) Fox, c) Huffington Post d) NYTimes
We observe that in Figure 3, the graphs for the articles that were predicted to be unbiased (a and c) both have visually sparse layouts. Meanwhile, the bias predicted articles (b and d) appear to be denser. Upon returning to the original data, we find that the news article from Figure 3c is actually labeled at the article level to be unbiased. Meanwhile, the original article label for Figure 3d is also labeled to be unbiased. Yet the article corresponding to graph 3d contains phrases like “Donald J. Trump leads a familiar refrain” and “Mr. Trump, who has been criticized for lacing his campaign speeches with profanity”, which some may consider to be inflammatory given word choice like “refrain” and “lacing” (Roappeport 2016). This example highlights the continued difficulty in developing reliable labels for an inherently subjective application. Furthermore, given the BASIL dataset only has three domains, we are not yet able to determine if these visual patterns are a spurious correlation.
To interrogate these patterns further, we look at the graphs of the articles using the article level “graph2vec_node_words” models from the VAX and OATH datasets as shown in Figures 4 and 5. As is apparent in the figures, there is no consistent visual pattern between the articles that are predicted to be unbiased and those that are predicted to be biased in either dataset. This reinforces our finding that a graph embedding approach is needed to capture the complex semantic structure characteristics of biased articles.
While these topic aligned datasets use weak bias labels, there is significantly greater variety in domains in these datasets in comparison to BASIL. This reduces the possibility the bias detection model trains on spurious domain specific characteristics. In validating our weak labels and looking into the article texts from Figures 4 and 5’s graphs, we find a clear usage of informational bias. The article corresponding to 5b ends with the sentence “Those words are directional instructions on Claymore anti-personal mines”(Brigham 2022), perhaps added to invoke a sense of incredulity of the actions of the Oath Keepers. Similarly, the AFN article from 4b opens with the title “Military docs drop vaccine bombshell but ‘fact-check’ has an explanation” (Jordahl and Billy 2022). The usage of “bombshell’ and quotation marks around “fact-check” are strong examples of lexical bias. Both articles align with their domain bias labels. On the other hand, the incorrectly classified OATH MSN article (Figure 5d), which is originally sourced from CNN, contains artifacts like “screaming match”, “The two continued to cut each other off…”, and “chastised”. Although not immediately clear without greater context which partisan lean it has, these informational and lexical rhetorical decisions lead to the assessment that this article is more biased than it is center. If the article had been correctly attributed to CNN in our dataset, it would have been aggregated as biased, rather than the unbiased, specifically left-center, label given to MSN (Rabinowitz 2016). Ultimately, this inspection demonstrates that while there are instances where our labeling assumptions falter, especially in borderline biased articles, our assumption on the generalizability of domain labels to weak article labels holds. Furthermore, our bias detection models may be more astute that these weak labels.
(a) Correct-Unbiased
(b) Correct-Biased
(c) Incorrect-True Biased
(d) Incorrect-True Unbiased
Figure 4: Graphs of articles with high probability predictions from the VAX dataset. Predictions from graph2vec_nodewords using binary labels (Macro F-1 = .75). Source domains are: a) ABC, b) AFN, c) Toronto Sun d) MSN
(a) Correct-Unbiased
(b) Correct-Biased
(c) Incorrect-True Biased
(d) Incorrect-True Unbiased
Figure 5: Graphs of articles with high probability predictions from the OATH dataset. Predictions from doc2vec_dbow_graph2vec_nodewords using binary labels (Macro F-1 = .76). Source domains are: a) MSN, b) Raw Story, c) Salon d) MSN
@@ 5 Conclusion and Future Work
Through our graph embeddings, we demonstrate that encoding semantic structure in news articles as co-occurrence networks can be used for both domain and article level bias detection. We couple unsupervised embeddings with a logistic regression to learn a separation of the embedding space. We also learn that biased texts have similar context agnostic semantic structures that can be captured by document networks (i.e. DocNet), regardless of the article topic and partisan lean. We show that you can get performance comparable to, and occasionally better than, transformer based language model by the simple inclusion of semantic structure in the form of word-co-occurrence networks. This is demonstrated by the success of the Graph2Vec embeddings in the supervised models, especially as it often has performance comparable to the SBERT embeddings, without the use of word embeddings. DocNet does not require pre-trained language models to generate word embeddings, nor does it require computationally expensive deep learning techniques. The most computationally time consuming portion of DocNet is the initial graph construction of an article. However, this can be parallelized across a corpus of documents for time savings. It also is an inductive approach, allowing for inexpensive and generalizable predictions on new datasets.
Through the varied groupings of bias labels, we identify that there is often insufficient information in the semantic structure of an article for accurate fine-grained bias classification. Future research using one-vs-rest classifiers rather than our generalized multinomial regression may be able to improve upon our results. This is an observation observed by other recent works (Carragher, Williams, and Carley 2024). On the other hand, fuzzier binary classifications can still be sufficiently characterized by semantic structures. This means that right-leaning news articles and left-leaning news articles have significantly more in common than either side of the political spectrum may believe. Although we are not able to pinpoint exactly what this writing style bias is due to the lack of interpretability in our graph embeddings, we can conclude that the characteristics of partisan rhetoric are more than context-dependent.
In future work, it would also be valuable to research embeddings that can fuse word embeddings with semantic structure. While our graph autoencoders perform better than a naive baseline, they are significantly less useful than our Graph2Vec embeddings in bias detection. Context represented in pre-trained language models is valuable to bias detection, as demonstrated in our SBERT model. Future developments in inductive graph embedding approaches that can efficiently and effectively represent dense networks and attributes would likely improve bias detection efforts. In the interim, our DocNet methodology provides a pathway for low-resourced bias detection, where computational resources, rigorously labeled datasets, or language models are not available. These findings may be especially advantageous to low resource languages and communities.